[figure: 9.7 cm]

STRONG HUBBARD CORRELATIONS AND WEAK MAGNETISM IN FeSi

N.Sluchanko^{a}, V.Glushkov^{a}, S.Demishev^{a}, L.Weckhuysen^{b}, V.Moshchalkov^{b}, A.Menovsky^{c} aGeneral Physics Institute of RAS, 38 Vavilov str., Moscow, 119991, Russia

bLaboratory Vaste-Stoffysica en Magnetisme, KULeuven, B-3001 Leuven, Belgium cVan der Waals-Zeeman Laboratory, 1018 XE Amsterdam, The Netherlands

The comprehensive study of a wide temperature range DC- and AC-transport and magnetization has been carried out in magnetic fields up to 120 kOe in the almost magnetic narrow-gap cubic semiconductor FeSi. The analysis of the experimental data obtained on high quality single crystals of iron monosilicide revealed that the Mott-Hubbard scenario of metal-insulator transition with on-site Coulomb interaction U≤3D (D is the band halfwidth) is likely to be the most appropriate description of the low temperature anomalies in transport, optical and thermodynamic characteristics in this strongly correlated electron system (N.E.Sluchanko et al., Phys. Rev. B 65 (2002), 064404). Temperature lowering induces density of states renormalization and formation of a short range (∼10Å) spin-polarons. This process is followed by the onset of coherent spin-fluctuations and transformation of the low carrier density (∼10^{17} cm^{−3}) heavy fermions into ferromagnetic microdomains at T_{c}≈15K. Detailed magnetic, magnetoresistance and microwave magnetoabsorption measurements allowed to deduce a low temperature magnetic T-H phase diagram of FeSi and estimate an exchange magnetic field in ferromagnetic microregions: H_{ex}≈350±100 kOe.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : N.E.Sluchanko

Address for correspondence : General Physics Institute of RAS, Vavilov str., 38, Moscow, 119991, Russia

Email address : nes@lt.gpi.ru